OpenStudy (anonymous):

Use linear interpolation and your line of fit for problem number 6 to estimate the number of human years which will correspond to 5 cat years. Cat Years 1 3 4 6 7 Human Years 13 33 43 63 73

OpenStudy (anonymous):

10*5+3=53 human years

OpenStudy (anonymous):

1. 10*1+3 = 13 3. 10*3+3 = 33 4. 10*4+3 = 43 6. 10*6+3 = 63 7. 10*7+3 = 73 so.. 5. 10*5+3 = 53